OpenStudy (anonymous):

The legs of a right triangle are 4 units and 7 units long. To the nearest whole number, what is the length of the hypotenuse?

OpenStudy (anonymous):

4^2+7^2=H^2

OpenStudy (anonymous):

so 16+49=H^2 H=\[\sqrt{65}\]

OpenStudy (anonymous):

Hashir is right. Using the Pythagorean Theorem, the answer you get is \[\sqrt{65}\]. But since the question asks for the nearest whole number, the answer should be: Hypotenuse = 8